### v1

# PO-Flow: Flow-based Generative Models for Sampling Potential Outcomes and Counterfactuals

Abstract
We propose PO-Flow, a novel continuous normalizing flow (CNF) framework for causal inference that jointly models potential outcomes and counterfactuals. Trained via flow matching, PO-Flow provides a unified framework for individualized potential outcome prediction, counterfactual predictions, and uncertainty-aware density learning. Among generative models, it is the first to enable density learning of potential outcomes without requiring explicit distributional assumptions (e.g., Gaussian mixtures), while also supporting counterfactual prediction conditioned on factual outcomes in general observational datasets. On benchmarks such as ACIC, IHDP, and IBM, it consistently outperforms prior methods across a range of causal inference tasks. Beyond that, PO-Flow succeeds in high-dimensional settings, including counterfactual image generation, demonstrating its broad applicability.
1 Introduction
Predicting potential outcomes (POs) lies at the core of causal inference. In a typical setting, each individual has two potential outcomes: $Y^{(1)}$ if they receive a treatment ($A=1$) and $Y^{(0)}$ if they do not ($A=0$). However, in the real world, only one outcome is observed, while the other potential outcome remains fundamentally unknown [28]. Accurately estimating both POs is crucial for decision-making in domains such as healthcare [41, 27, 50], where clinicians may wish to compare a patient’s likely outcomes under different treatment options [60, 20].
In the causal inference literature, the vast majority of methods focus on estimating the average treatment effect (ATE) and the conditional average treatment effect (CATE) at the population level [32, 25, 1, 22, 37, 42, 55, 67]. It is important to note that although some methods perform well in estimating population-level CATE, they may underperform in individualized potential outcome (PO) prediction, as accurately predicting POs in finite samples is challenging [52].
However, making decisions for individuals based on average treatment effects can be misleading. Therefore, accurately predicting individualized POs and learning their full distributions with uncertainty is a crucial task [69, 35]. Recent works [67, 44, 45] have explored generative models to sample POs and estimate empirical distributions. However, methods that learn the full density of POs without relying on explicit distributional assumptions (e.g., Gaussian mixtures) remain scarce.
In addition, another critical challenge is estimating the counterfactual outcome conditioned on the factual—that is, predicting what would have happened under the alternative treatment, given the observed outcome. This task is especially relevant in post-hoc clinical analyses, where practitioners ask, “Would the patient have fared better on another therapy?” [66, 29, 21].
Counterfactual prediction differs from standard PO estimation—it requires reasoning about outcomes under an alternative treatment conditioned on the observed outcome, introducing dependencies between observed and unobserved outcomes. Structural Causal Models (SCMs) [48] and modern extensions [46, 6, 13, 7, 65, 49, 9] offer the most widely used theoretical framework for counterfactual reasoning. SCMs methods typically assume known causal structures or aim to learn them explicitly. Yet, it may be possible to estimate counterfactuals directly without recovering the SCM [70]. In PO-Flow, we do not explicitly assume an SCM, but seek to directly estimate counterfactual outcomes.
We aim to develop a generative modeling framework that jointly supports population-level CATE estimation, individualized PO prediction with density modeling, and counterfactual outcome prediction without explicitly modeling an SCM—an area that has been relatively less explored.
1.1 Related Works
CATE Estimation: Numerous methods have been proposed for CATE estimation [32, 25, 1, 22, 37, 42, 55, 67]. Broadly, these fall into model-agnostic and model-specific approaches. Model-agnostic estimators, or meta-learners [37, 15, 16, 2, 54], typically follow one of two strategies [43]: (a) one-step methods, which fit regression models and compute CATE as the difference in predicted outcomes, and (b) two-step methods, which estimate nuisance functions to construct pseudo-outcomes, then regress them on covariates.
In contrast, model-specific estimators adapt machine learning techniques for CATE estimation, including methods such as [14, 55, 42, 34, 67]. More recently, neural networks have been widely applied to this task, with notable examples in [55, 42, 67]. However, most of these approaches mainly focus on estimating CATE rather than predicting individualized potential outcomes.
Individualized POs Prediction: Although some CATE methods can be adapted for potential outcomes (POs) prediction, they may perform suboptimally, particularly on finite samples, as they are not specifically optimized for accurate PO estimation. This limitation partly arises because many methods assume that CATE has a simpler functional structure than POs. Nevertheless, several CATE approaches are also designed with POs prediction as the target [37, 55, 67, 42, 45, 43], and these will serve as our primary baselines for experimental comparisons.
Counterfactual Outcomes Prediction: Structural Causal Models (SCMs) provide a well-established framework for counterfactual reasoning [48, 46, 6, 13, 7, 65, 49, 9], but practical applications often require specifying the full causal graph and structural equations, which may be difficult or infeasible for real-world data. Some recent works avoid explicitly modeling SCMs [39, 5, 44, 62], but typically address specific tasks and rely on additional information, such as temporal structure.
Some works [27, 34, 42, 38, 67] targeting CATE estimation describe their goal as counterfactual inference. However, in these works, "counterfactual" typically refers to the unobserved potential outcome under an alternative treatment, without conditioning on the observed factual outcome.
While SCMs offer theory-grounded counterfactual reasoning, general-purpose methods for individualized counterfactual prediction—without explicitly modeling SCMs or relying on auxiliary information—remain scarce. Although [67, 43] were not originally designed for counterfactuals, we adapt them by incorporating factual information into intermediate steps and use them as baselines.
Diffusion and Flow Models: Diffusion models [26, 57, 58] and flow-based generative models [12, 64, 40, 4] have demonstrated strong performance across a range of applications, including image generation, statistical sampling [18, 59, 61], and physics-informed modeling [8, 68, 51].
Recently, diffusion and flow-based generative models have been applied to causal inference tasks such as counterfactual image generation [53, 19], causal query answering [10], and inference over causal graphs or structural causal models [9, 33].
However, to our knowledge, only two generative algorithms have been proposed specifically for the potential outcomes (PO) task: the diffusion-based DiffPO [43] and the discrete normalizing flow-based INFs [45]. We conduct extensive comparisons with both methods in this work.
1.2 Main Contributions
We propose PO-Flow, a continuous normalizing flow (CNF) framework for potential outcomes and counterfactuals. Our main contributions are: (1) PO-Flow provides a unified generative framework for potential outcome prediction, CATE estimation, and general-purpose counterfactual prediction—without explicitly modeling structural causal models (SCMs) or relying on auxiliary information for the counterfactual task; (2) PO-Flow enables learning the density of predicted potential outcomes, offering deeper insights into uncertainty; (3) PO-Flow outperforms state-of-the-art methods across a range of causal tasks on multiple benchmarks.
2 Preliminaries
2.1 Neural ODE and Continuous Normalizing Flow:
A Neural ODE is a continuous model where the trajectory of data is modeled as the solution of an ordinary differential equation (ODE). Formally, in $\mathbb{R}^{d}$, given an input $y_{0}=y(0)$ at time $t=0$, the transformation to the output $y_{1}=y(1)$ at time $t=1$ is governed by:
$\frac{dy(t)}{dt}=v_{\theta}(y(t),t),\quad t\in[0,1],$ (1)
where $v_{\theta}:\mathbb{R}^{d_{Y}+1}\to\mathbb{R}^{d_{Y}}$ represents the velocity field, which is parameterized by a neural network with parameter $\theta$. The time horizon is rescaled to $[0,1]$ without loss of generality.
Continuous Normalizing Flow (CNF) is a class of normalizing flows in which the transformation of a probability density is governed by a time-continuous Neural ODE. The marginal density of $y(t)$, denoted as $p(y,t)$, evolves according to the continuity equation implied by the ODE in Eq. (1). This continuity equation is given by:
$\partial_{t}p(y,t)+\nabla\cdot(p(y,t)v_{\theta}(y,t))=0,\quad t\in[0,1],$ (2)
[…]
When the Neural ODE is well-posed, it defines a continuous and invertible mapping from the initial sample $y_{0}$ to the terminal sample $y_{1}$. The inverse mapping from $y_{1}$ to $y_{0}$ can be computed by integrating the ODE (1) in reverse time. This framework allows setting $y_{0}\sim p(\cdot,0)$ as the data distribution and $y_{1}\sim p(\cdot,1)$ as the base (noise) distribution, typically chosen as $N(0,I)$.
For notational convenience, we define $p_{Y}(\cdot):=p(\cdot,0)$ as the data distribution and $q(\cdot):=p(\cdot,1)$ as the base (noise) distribution. Throughout this paper, we adopt the common choice $q(\cdot)=N(0,I)$.
2.2 Flow Matching
Flow Matching (FM) [40, 4] provides a simulation-free approach for training continuous normalizing flows by directly regressing the model velocity $v_{\theta}(y,t)$ onto a predefined target velocity field $u(y,t)$. The original FM objective is defined as:
$\mathcal{L}_{\text{FM}}(\theta)=\mathbb{E}_{t\sim\mathcal{U}[0,1],y\sim p(\cdot,t)}\left[\|v_{\theta}(y,t)-u(y,t)\|^{2}\right],$ (3)
where $u(y,t)$ is an analytically specified target velocity field.
Linear Interpolant: Due to the intractability of $u(y,t)$ without conditioning on the initial data $y_{0}$, Conditional Flow Matching (CFM) is introduced. Conditioned on an observed sample $y_{0}\sim p_{Y}(\cdot)$, CFM employs a pre-specified interpolation function $I_{t}$ between two endpoints $y_{0}$ and $y_{1}\sim q(\cdot)$, where $q$ is typically $N(0,I)$. The interpolated path is defined as:
$\phi_{t}:=I_{t}(y_{0},y_{1}),\quad t\in[0,1].$ (4)
Here, $I_{t}$ can be analytically specified. Conditioned on $y_{0}$, the optimal and most information-preserving choice is the linear interpolation path [40]:
$\phi_{t}=(1-t)y_{0}+(t+\sigma_{\text{min}}(1-t))y_{1},$ (5)
where $y_{1}\sim N(0,I)$ and $\sigma_{\text{min}}$ is a small positive hyperparameter that ensures $p(\phi_{t},0)\sim N(y_{0},\sigma_{\text{min}}^{2}I)$. When $\sigma_{\text{min}}=0$, this reduces to the strict linear interpolant $\phi_{t}=(1-t)y_{0}+ty_{1}$.
Here, $\phi_{t}$ serves as a pre-specified reference trajectory that the ODE solution $y(t)$ should follow at time $t$. Under the linear interpolant defined in (5), the corresponding reference velocity field becomes:
$\frac{d\phi_{t}}{dt}=(1-\sigma_{\text{min}})y_{1}-y_{0}.$ (6)
Training loss: CFM trains the flow $v_{\theta}(y(t),t)$ by directly regressing it onto the reference velocity field (6). The training objective is given by:
$\mathcal{L}_{\text{CFM}}(\theta)=\ \mathbb{E}_{t\sim\mathcal{U}[0,1],y_{0}\sim p_{Y}(\cdot),y_{1}\sim q(\cdot)}\|v_{\theta}(\phi_{t},t)-\frac{d\phi_{t}}{dt}\|^{2}.$ (7)
3 Problem Settings
Notations and Setups: We consider a setting where each individual is described by covariates $X\in\mathbb{R}^{d_{X}}$, a treatment assignment $A\in\{0,1\}$, and an outcome $Y^{(A)}\in\mathbb{R}^{d_{Y}}$ under treatment $A$. The observational dataset is $\mathcal{D}=\{(y^{(a)}_{i},x_{i},a_{i})\}_{i=1}^{n}\sim p_{Y,X,A}(\cdot)$, where $y_{i}^{(a)}$ is the observed potential outcome under treatment $a_{i}$, and $(x_{i},a_{i})\sim p_{X,A}(\cdot)$ are the associated covariates and treatment.
Potential Outcomes: In the potential outcomes framework, each individual has two potential outcomes: $Y^{(0)}$ (control) and $Y^{(1)}$ (treatment), but only one is observed based on assignment $A$. We define potential outcome prediction as estimating $Y^{(a)}\mid X,A=a$. For convenience, we denote $Y^{(a)}$ and $\hat{Y}^{(a)}$ as the true and predicted potential outcomes.
Counterfactual Outcomes: In contrast to potential outcomes, we define the counterfactual outcome as $Y_{\text{cf}}^{(1-a)}:=Y^{(1-a)}\mid X,A=a,Y^{(a)}$: the outcome had the individual received the opposite treatment, conditioned on its observed potential (factual) outcome. For convenience, we denote $Y_{\text{cf}}^{(1-a)}$ and $\hat{Y}_{\text{cf}}^{(1-a)}$ as the true and predicted counterfactual outcomes, respectively.
To ensure the identifiability of potential outcomes, we adopt the standard assumptions commonly used in the literature [43, 45, 67]:
Assumption 3.1.
(1) Consistency: If an individual receives treatment $A=a$, the observed outcome equals the corresponding potential outcome: $Y=Y^{(a)}$. (2) Unconfoundedness: All confounders are observed, i.e., $\{Y^{(0)},Y^{(1)}\}\perp\!\!\!\perp A\mid X$. (3) Overlap: Treatment assignment is probabilistic: $10<P(A=1\mid X=x)<$ 1, for all $x\in\mathcal{X}$.
Counterfactuals are rarely identifiable without strong assumptions [63, 47]. Estimating them can often be framed as an optimal transport problem [11, 17]. To constrain the space of joint transport mappings, our PO-Flow framework introduces the following assumption, applicable only to Section 4.2:
Assumption 3.2.
There exists a deterministic and invertible mapping $f:\mathbb{R}^{d_{Y}+d_{X}}\times\{0,1\}\rightarrow\mathbb{R}^{d_{Z}}$ that encodes the observed outcome as $Z_{Y}=f(Y^{(A)},X,A)$. A counterfactual latent representation can be obtained as $Z^{\prime}_{Y}=Z_{Y}+\delta$, where $\delta\sim\mathcal{N}(0,\Sigma(X))$ is an optional covariate-dependent noise term. The counterfactual is then recovered via $Y_{\text{cf}}^{(1-A)}=f^{-1}(Z^{\prime}_{Y},X,1-A)$.
4 PO-Flow: A Continuous Normalizing Flow for Potential Outcomes
We introduce PO-Flow, a continuous normalizing flow framework for modeling potential and counterfactual outcomes, with three main goals: 1. predicting individualized potential outcomes, 2. estimating counterfactual outcomes, and 3. learning the densities.
Our PO-Flow framework is built on Continuous Normalizing Flows (CNFs) and is trained using the Flow Matching (FM) objective. Our learning target is the trajectory of the sample $y(t)$ over $t\in[0,1]$. To condition the dynamics on both covariates $x$ and treatment $a$, our Neural ODE is defined as:
$\small\frac{dy(t)}{dt}=v_{\theta}(y(t),t,x,a),\quad t\in[0,1],$ (8)
where $v_{\theta}:\mathbb{R}^{d_{Y}+d_{X}+2}\rightarrow\mathbb{R}^{d_{Y}}$ is parametrized by a neural network.
Notation: Given a data point $(y^{(a)},x,a)\sim p_{Y,X,A}(\cdot)$, we define $y_{0}=y^{(a)}$ as the observed training sample at time $t=0$ in the Neural ODE, and $y_{1}\sim q(\cdot)$ as the noise sample at time $t=1$. Recall that we denote $\hat{y}^{(a)}$ and $\hat{y}_{\text{cf}}^{(1-a)}$ as the predicted potential and counterfactual outcomes, respectively.
Training loss: PO-Flow adopts the Conditional Flow Matching (CFM) objective for training. Under the updated definitions, our training loss retains the form of Equation (7) and is given by:
$\min_{\theta}\ \mathcal{L}_{\text{CFM}}(\theta)=\ \mathbb{E}_{t\sim\mathcal{U}[0,1],p_{Y,X,A}(\cdot),q(\cdot)}\|v_{\theta}(\phi_{t},t,x,a)-\frac{d\phi_{t}}{dt}\|^{2},$ (9)
where $\phi_{t}$ and $\frac{d\phi_{t}}{dt}$ are defined in (5) and (6), respectively.
Remark: Following prior work [23, 43], one may estimate propensity scores $w(x)=p(A=1\mid X=x)$ using a pre-trained classifier and apply inverse probability weighting $\frac{1}{w(x)}$ in the loss. However, we found this did not improve performance, but both implementations are included in our code.
Forward Process: Given the velocity field $v_{\theta}$, the forward process refers to pushing a factual outcome $y^{(a)}$ (at $t=0$) to a latent embedding $z_{y}$ (at $t=1$) via:
[…]
Through the forward process, $z_{y}$ encodes information from the factual outcome $y^{(a)}$ as well as its associated covariates $x$ and treatment $a$.
Reverse Process: The reverse process is initialized differently for interventional and counterfactual outcome prediction, as detailed in Sections 4.1 and 4.2. In general, given a noise sample $y_{1}\sim q(\cdot)$ at $t=1$, the reverse process is defined as:
[…]
Empirically, we use the Runge-Kutta method for numerical integration, with implementation details provided in Appendix B.4.
During training, we construct synthetic paths $\phi_{t}$ between $y_{0}$ and $y_{1}\sim q(\cdot)$ using the CFM loss, conditioning on $y_{0}$. However, Proposition 4.1 shows that the gradient of the CFM loss with respect to the model parameters is equivalent to the gradient of the original FM loss.
Proposition 4.1.
Assume that $p_{t}(y|x,a)>0$ for all $y\in\mathbb{R}^{d_{Y}}$ and $t\in[0,1]$. Then, up to a constant independent of $\theta$, the Conditional Flow Matching (CFM) loss and the original Flow Matching (FM) loss are equivalent. Hence,
$\nabla_{\theta}\mathcal{L}_{\text{FM}}(\theta)=\nabla_{\theta}\mathcal{L}_{\text{CFM}}(\theta).$
As a result of Proposition 4.1, the learned velocity field $v_{\theta}$ during the reverse process generates samples $\hat{y}_{0}$ independent of the individual samples $y_{0}$ used during training. Consequently, since $y_{1}\sim q(\cdot)$ is random, $v_{\theta}$ induces a distribution over outcomes conditioned on $(x,a)$.
4.1 PO-Flow for Potential Outcomes Prediction
  Require: Trained $v_{\theta}$
  Input: $(x,a)\sim p_{X,A}(\cdot)$
  1: $y_{1}\sim q(\cdot)$
[…]
  Return $\hat{y}^{(a)}:=\hat{y}_{0}$
Potential Outcomes Prediction
Algorithm 1
Potential Outcomes Prediction
The potential outcomes prediction in PO-Flow follows a simple reverse process, as defined in Equation (11). Starting from a noise sample $y_{1}\sim q(\cdot)$, we integrate the velocity field $v_{\theta}$ backward in time, conditioned on covariates $x$ and treatment $a$, to obtain a point estimate of the potential outcome $\hat{y}^{(a)}$. The procedure is summarized in Algorithm 1.
In addition to predicting point estimates of potential outcomes, PO-Flow is capable of learning the probability densities of the potential outcomes, as detailed in Section 4.3.
4.2 PO-Flow for Counterfactual Outcomes Prediction
  Require: Trained $v_{\theta}$
  Input: $(y^{(a)},x,a)\sim p_{Y,X,A}(\cdot)$
  1. Forward process (encoding):
[…]
         with $y(0)=y^{(a)}$.
  3. Uncertainty-aware perturbations:
  4.     $\delta\sim N(0,\Sigma(x))$
  5.     $z^{\prime}_{y}=z_{y}+\delta$
  6. Reverse process (decoding):
[…]
         with $y(1)=z^{\prime}_{y}$.
  Return: $\hat{y}_{\text{cf}}^{(1-a)}:=\hat{y}_{0}$
Counterfactual Prediction
Algorithm 2
Counterfactual Prediction
Under Assumption 3.2, the counterfactual outcome prediction in PO-Flow consists of both a forward and a reverse process. Given a factual observation $(y^{(a)},x,a)\sim p_{Y,X,A}(\cdot)$, the forward process (Lines 2–3 in Algorithm 2) encodes the information $(y^{(a)},x,a)$ into a latent representation centered at $z_{y}$. Lines 4–6 introduce uncertainty-aware perturbations to $z_{y}$, reflecting covariate-dependent uncertainty. Finally, the reverse process (Lines 7–8) decodes the latent $z_{y}^{\prime}$ back into the counterfactual outcome prediction $\hat{y}_{\text{cf}}^{(1-a)}$.
Figure 1: Illustration of counterfactual predictions. Left: two potential outcomes distribution. Right: Base (noise) distribution.
Because of the smooth and information-preserving nature of the PO-Flow path trained via flow matching, the encoded $z_{y}$ retains rich information from the factual outcome $y^{(a)}$, enabling accurate decoding of the counterfactual $\hat{y}_{\text{cf}}^{(1-a)}$. To our knowledge, this is the first work to explore counterfactual outcome prediction using flow or diffusion-based models.
As illustrated in Figure 1, diffusion models rely on noisy and stochastic forward processes, which degrade the information in $z_{y}$ and lead to reversed samples that resemble interventional potential outcomes rather than counterfactuals. In Section 5.4, we present counterfactuals using an adapted implementation of DiffPO [43] (originally not designed for counterfactual), compared with PO-Flow.
4.3 PO-Flow for Learning Potential Outcomes Densities
Another distinctive advantage of PO-Flow is its ability to learn the density of the Potential Outcomes (POs). We define $\hat{p}_{Y}(\cdot\mid x,a,y_{1})$ as the PO density learned by PO-Flow, conditioned on $(x,a)$ and a base sample $y_{1}$. The explicit form of the learned PO density is presented in Proposition 4.2, with the proof provided in Appendix A.
Proposition 4.2.
Given a noise sample $y_{1}\sim q(\cdot)$ at $t=1$, the log-density of the resulting potential outcome at $t=0$, obtained via PO-Flow, is given by:
$\log\hat{p}_{Y}(\cdot|x,a,y_{1})=\log q(y_{1})+\int_{0}^{1}\nabla\cdot v_{\theta}(y(t),t,x,a)ds.$ (12)
Empirically, the divergence of the velocity field can be computed under the Hutchinson trace estimator (e.g., [31, 64]). More details are provided in Appendix B.5.
$\small\nabla\cdot v_{\theta}(y(t),t,x,a)\approx\mathbb{E}_{p(\epsilon)}\left[\epsilon\cdot\frac{v_{\theta}(y(t)+\sigma\epsilon,t,x,a)-v_{\theta}(y(t),t,x,a)}{\sigma}\right],$ (13)
where $p(\epsilon)$ is a distribution in $\mathbb{R}^{d_{Y}}$ satisfying $\mathbb{E}[\epsilon]=0$ and Cov$(\epsilon)=I$ (e.g., a standard Gaussian).
To our knowledge, PO-Flow is the first generative approach in causal inference to explicitly learn the probability density of predicted potential outcomes. Unlike prior works [44, 45] that generate PO samples and estimate empirical distributions, PO-Flow provides direct access to likelihoods, enabling deeper insights into uncertainty.
This density estimation allows PO-Flow to assess prediction confidence via log-likelihood, identify unlikely outcomes, and quantify uncertainty. For counterfactual prediction, one can perturb the latent code $z_{y}^{\prime}$ multiple times and select the most confident $\hat{y}_{\text{cf}}^{(1-a)}$ based on log-likelihood, effectively reducing variance while improving robustness.
5 Experiments
Baselines: Our baselines cover several categories: model-agnostic learners that learn the conditional distribution means (S-learner and T-learner [37]), representation-based conditional regressors (Counterfactual Regression (CFR) [55]), generative approaches (Causal Effect Variational Autoencoder (CEVAE) [42]; GAN for Individualized Treatment Effect (GANITE) [67]), and most recent generative approaches based on diffusion (DiffPO [43]) and discrete normalizing flows (INFs [45]).
Datasets: Counterfactual outcomes and unobserved potential outcomes (POs) are inaccessible in real-world data but can be simulated in semi-synthetic settings, as discussed in detail in Appendix B.1. Accordingly, we evaluate PO-Flow on standard synthetic benchmarks, including ACIC 2016, ACIC 2018, IHDP, and IBM.
The code and datasets are available at https://github.com/StatFusion/PO-Flow.
5.1 Prediction of Potential Outcomes
Table 1: Root mean squared error (RMSE) of estimated potential outcomes (POs) across different methods on the ACIC 2018, IHDP, and IBM datasets. Results are averaged over 10-fold cross-validation.
ACIC 2018 IHDP IBM
$\text{RMSE}_{\text{in}}$ $\text{RMSE}_{\text{out}}$ $\text{RMSE}_{\text{in}}$ $\text{RMSE}_{\text{out}}$ $\text{RMSE}_{\text{in}}$ $\text{RMSE}_{\text{out}}$
[…]
DiffPO $0.72_{\pm.14}$ $0.79_{\pm.16}$ $1.33_{\pm.20}$ $1.40_{\pm.22}$ $1.82_{\pm.38}$ $1.85_{\pm.35}$
INFs $0.69_{\pm.16}$ $0.78_{\pm.17}$ $1.02_{\pm.10}$ $\mathbf{1.20_{\pm.12}}$ $1.52_{\pm.24}$ $1.57_{\pm.28}$
S-learner $\mathbf{0.54_{\pm.09}}$ $\mathbf{0.57_{\pm.11}}$ $1.31_{\pm.18}$ $1.44_{\pm.20}$ $1.01_{\pm.17}$ $1.16_{\pm.19}$
T-learner $1.57_{\pm.32}$ $1.71_{\pm.40}$ $1.45_{\pm.25}$ $1.49_{\pm.27}$ $1.89_{\pm.47}$ $1.96_{\pm.50}$
CEVAE $0.83_{\pm.17}$ $0.85_{\pm.17}$ $1.18_{\pm.15}$ $1.36_{\pm.18}$ $0.96_{\pm.12}$ $\mathbf{0.98_{\pm.12}}$
CFR $0.94_{\pm.10}$ $0.98_{\pm.12}$ $1.10_{\pm{.20}}$ $\mathbf{1.18_{\pm{.20}}}$ $2.09_{\pm.45}$ $2.17_{\pm.48}$
GANITE $0.88_{\pm.15}$ $0.97_{\pm.15}$ $1.60_{\pm.36}$ $1.67_{\pm.34}$ $2.48_{\pm.39}$ $2.59_{\pm.47}$
Evaluation Metrics: We evaluate the point estimates of potential outcomes (POs) using the root mean squared error (RMSE) between the estimated outcomes $\hat{y}_{i}$ and the true outcomes $y_{i}$, computed as $\text{RMSE}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}}$. $\text{RMSE}_{\text{in}}$ and $\text{RMSE}_{\text{out}}$ refer to the RMSE evaluated on the training and test sets, respectively.
Results: Table 1 shows that PO-Flow achieves the most accurate prediction of potential outcomes on the majority of datasets.
5.2 Estimation of CATEs
Conditional Average Treatment Effect (CATE): The CATE is defined as $\tau(x)=\mathbb{E}[y_{1}-y_{0}\mid X=x]$, representing the expected treatment effect conditioned on covariates $x$. It captures individual-level heterogeneity in causal effects.
Evaluation Metrics: We evaluate CATE using the precision in estimation of heterogeneous effect (PEHE), defined as the mean squared error between estimated and true effects: $\epsilon_{\text{PEHE}}=\frac{1}{n}\sum_{i=1}^{n}(\hat{\tau}(x_{i})-\tau(x_{i}))^{2}$. Table 5 reports the root PEHE for both training and test sets. PO-Flow achieves the most accurate CATE estimates on most datasets.
Table 2: Root Precision in Estimation of Heterogeneous Effect ($\sqrt{\epsilon_{\text{PEHE}}}$) across different methods on the ACIC 2016, ACIC 2018, IHDP, and IBM datasets. Results are averaged over 10-fold cross-validation.
ACIC 2016 ACIC 2018 IHDP IBM
$\sqrt{\epsilon_{\text{PEHE}}^{\text{in}}}$ $\sqrt{\epsilon_{\text{PEHE}}^{\text{out}}}$ $\sqrt{\epsilon_{\text{PEHE}}^{\text{in}}}$ $\sqrt{\epsilon_{\text{PEHE}}^{\text{out}}}$ $\sqrt{\epsilon_{\text{PEHE}}^{\text{in}}}$ $\sqrt{\epsilon_{\text{PEHE}}^{\text{out}}}$ $\sqrt{\epsilon_{\text{PEHE}}^{\text{in}}}$ $\sqrt{\epsilon_{\text{PEHE}}^{\text{out}}}$
[…]
DiffPO $0.91_{\pm.20}$ $1.12_{\pm.26}$ $0.07_{\pm.03}$ $\mathbf{0.08_{\pm.04}}$ $0.79_{\pm.16}$ $0.84_{\pm.18}$ $0.08_{\pm.03}$ $0.09_{\pm.03}$
INFs $1.37_{\pm.28}$ $1.42_{\pm.30}$ $0.33_{\pm.11}$ $0.37_{\pm.12}$ $0.90_{\pm.27}$ $0.95_{\pm.27}$ $0.16_{\pm.08}$ $0.18_{\pm.08}$
S-learner $3.81_{\pm.76}$ $3.87_{\pm.98}$ $\mathbf{0.06_{\pm.02}}$ $\mathbf{0.07_{\pm.03}}$ $0.92_{\pm.07}$ $0.93_{\pm.08}$ $0.06_{\pm.02}$ $0.07_{\pm.04}$
T-learner $5.02_{\pm.71}$ $5.15_{\pm.87}$ $1.21_{\pm.47}$ $1.32_{\pm.53}$ $1.24_{\pm.08}$ $1.31_{\pm.09}$ $0.60_{\pm.14}$ $0.69_{\pm.15}$
CEVAE $2.06_{\pm.61}$ $2.31_{\pm.74}$ $1.19_{\pm.15}$ $1.21_{\pm.16}$ $1.41_{\pm.10}$ $1.43_{\pm.10}$ $0.73_{\pm.15}$ $0.75_{\pm.16}$
CFR $2.80_{\pm.58}$ $2.74_{\pm.60}$ $0.77_{\pm.18}$ $0.80_{\pm.18}$ $0.73_{\pm.04}$ $0.78_{\pm.04}$ $1.05_{\pm.27}$ $1.08_{\pm.28}$
GANITE $3.02_{\pm.76}$ $3.11_{\pm.74}$ $0.55_{\pm.11}$ $0.60_{\pm.13}$ $1.90_{\pm.40}$ $2.40_{\pm.40}$ $0.77_{\pm.20}$ $0.80_{\pm.24}$
5.3 Learning the Full Potential Outcome Distributions
Beyond point estimation, PO-Flow offers two distinct advantages for modeling potential outcomes: (1) it enables individualized sampling from the conditional distribution $P(Y\mid X,A)$, generating a rich set of plausible outcomes for each individual; and (2) it supports learning the probability density of any generated outcome, providing deeper insights into predictive uncertainty.
5.3.1 Learning potential outcome densities
PO-Flow enables direct computation of potential outcome densities, as shown in Proposition 4.2. To our knowledge, it is the first generative method in causal potential outcomes to provide this capability.
Evaluation Metrics: To evaluate the learned densities, we compute the KL divergence between the PO-Flow estimate $\hat{p}(\cdot|x,a)$ and the ground-truth density $p_{\text{true}}(\cdot|x,a)$. Given that the semi-synthetic datasets assume $y^{(a)}\sim\mathcal{N}(\mu(a),1)$, the individual-level KL is $\text{KL}(\hat{Y}_{X}^{(a)}\|Y_{X}^{(a)})=\mathbb{E}_{\hat{y}\sim\hat{p}}\left[\log\frac{\hat{p}(\hat{y}|x,a)}{\mathcal{N}(\mu(\hat{y}),1)}\right]$, where $\hat{p}$ is computed via (12). Table 3 reports the average results.
Table 3: KL divergence of learned potential outcome distributions across datasets. Results are averaged over 10-fold cross-validation.
ACIC 2016 ACIC 2018 IHDP IBM
PO-Flow $\mathbf{0.14_{\pm.04}}$ $\mathbf{0.08_{\pm.02}}$ $\mathbf{0.09_{\pm.03}}$ $\mathbf{0.08_{\pm.03}}$
GP-ITE $1.73_{\pm.34}$ $0.73_{\pm.19}$ $1.42_{\pm.30}$ $0.68_{\pm.17}$
Baselines: A natural baseline with density estimation capability is Gaussian Processes for interventional treatment effects, as used in [3, 30].
Results: PO-Flow achieves much lower KL divergence than GP baselines on the ACIC 2018, IHDP, and IBM datasets, indicating more accurate density estimation of potential outcomes.
This property allows PO-Flow to assess prediction confidence via log-density and to identify unlikely outcomes. For counterfactual estimation, one can sample multiple candidates from a perturbed latent region and select the most likely one based on the model’s estimated density.
5.3.2 Empirical distribution estimations
Evaluation Metrics: To enable comparison of estimated potential outcome distributions across methods, we include empirical distribution estimates for all baselines. These are evaluated using the Wasserstein-1 distance, defined as: $W_{1}(\hat{p}(y|x,a),p(y|x,a))=\inf_{\gamma\in\Gamma(\hat{p},p)}\mathbb{E}_{(\hat{y},y\sim\gamma)}[\|\hat{y}-y\|_{1}]$. The results are shown in Table 4.
Table 4: In-sample and out-of-sample Wasserstein distances across different methods on the ACIC 2016, ACIC 2018, IHDP, and IBM datasets. Results are averaged over 10-fold cross-validation.
ACIC 2016 ACIC 2018 IHDP IBM
$\hat{W}_{1}^{\text{in}}$ $\hat{W}_{1}^{\text{out}}$ $\hat{W}_{1}^{\text{in}}$ $\hat{W}_{1}^{\text{out}}$ $\hat{W}_{1}^{\text{in}}$ $\hat{W}_{1}^{\text{out}}$ $\hat{W}_{1}^{\text{in}}$ $\hat{W}_{1}^{\text{out}}$
[…]
DiffPO $0.47_{\pm.17}$ $0.51_{\pm.18}$ $\mathbf{0.04_{\pm.02}}$ $\mathbf{0.13_{\pm.05}}$ $0.80_{\pm.15}$ $0.88_{\pm.17}$ $0.69_{\pm.14}$ $0.75_{\pm.14}$
INFs $0.59_{\pm.13}$ $0.67_{\pm.14}$ $0.08_{\pm.03}$ $0.16_{\pm.05}$ $0.32_{\pm.08}$ $\mathbf{0.40_{\pm.09}}$ $0.36_{\pm.08}$ $\mathbf{0.45_{\pm.09}}$
S-learner $1.01_{\pm.13}$ $1.02_{\pm.20}$ $0.35_{\pm.13}$ $0.49_{\pm.20}$ $0.71_{\pm.14}$ $0.85_{\pm.19}$ $0.46_{\pm.11}$ $0.55_{\pm.14}$
T-learner $1.22_{\pm.15}$ $1.48_{\pm.18}$ $1.04_{\pm.13}$ $1.00_{\pm.30}$ $0.83_{\pm.20}$ $0.92_{\pm.28}$ $1.21_{\pm.23}$ $1.67_{\pm.32}$
CEVAE $0.62_{\pm.17}$ $0.70_{\pm.21}$ $0.88_{\pm.18}$ $1.02_{\pm.23}$ $0.65_{\pm.14}$ $0.76_{\pm.20}$ $0.38_{\pm.07}$ $\mathbf{0.45_{\pm.07}}$
CFR $0.68_{\pm.13}$ $0.73_{\pm.15}$ $0.88_{\pm.12}$ $0.90_{\pm.15}$ $0.50_{\pm.17}$ $0.56_{\pm.19}$ $0.92_{\pm.20}$ $0.93_{\pm.21}$
GANITE $0.74_{\pm.20}$ $0.86_{\pm.23}$ $0.89_{\pm.24}$ $1.18_{\pm.25}$ $0.71_{\pm.23}$ $0.82_{\pm.29}$ $1.09_{\pm.27}$ $1.34_{\pm.30}$
Figure 2: Comparison between empirically estimated and the ground-truth POs distributions. Left: $P(y|x,a=0)$ on IHDP. Right: $P(y|x,a=1)$ on IHDP.
Understanding the distribution of potential outcomes (POs) is essential for reasoning about uncertainty in treatment decisions. As shown in Figure 2, PO distributions can shift differently across treatments for the same individual. PO-Flow provides not only point estimates but also distributional insight, supporting uncertainty-aware analysis. Section 5.3.1 further demonstrates its ability to estimate densities accurately. This enables practitioners to assess whether an outcome lies within a desired range and evaluate its likelihood (Proposition 4.2) under a given treatment, offering a clearer sense of confidence in the effect.
5.4 Counterfactual Outcome Estimation
Evaluation Metrics: We evaluate the predicted counterfactual outcomes using the root mean squared error (RMSE) between the estimated counterfactuals $\hat{y}_{i,\text{cf}}^{(1-a)}$ and the ground-truth counterfactuals $y_{i,\text{cf}}^{(1-a)}$, computed as $\text{RMSE}=\small\sqrt{\frac{1}{n}\sum_{i=1}^{n}(y_{i,\text{cf}}^{(1-a)}-\hat{y}_{i,\text{cf}}^{(1-a)})^{2}}$. The results are presented in Table 5. Here, $\text{RMSE}_{\text{in}}$ and $\text{RMSE}_{\text{out}}$ denote the RMSE evaluated on the training and test sets, respectively.
Baselines: To our knowledge, very few existing methods are specifically designed for counterfactual outcome prediction conditioned on factuals in general settings—without relying on structural causal models or temporal information. While methods like CFR include “counterfactual” in name, they target unobserved potential outcomes rather than factual-conditioned ones, though we still include them as valuable baselines. We also adapt DiffPO and GANITE for this task, as detailed in Appendix B.6.
Table 5: Root mean squared error (RMSE) of estimated counterfactual outcomes across different methods on the ACIC 2016, ACIC 2018, IHDP, and IBM datasets. Results are averaged over 10-fold cross-validation.
ACIC 2016 ACIC 2018 IHDP IBM
$\text{RMSE}_{\text{in}}$ $\text{RMSE}_{\text{out}}$ $\text{RMSE}_{\text{in}}$ $\text{RMSE}_{\text{out}}$ $\text{RMSE}_{\text{in}}$ $\text{RMSE}_{\text{out}}$ $\text{RMSE}_{\text{in}}$ $\text{RMSE}_{\text{out}}$
[…]
DiffPO $1.77_{\pm.20}$ $1.78_{\pm.20}$ $0.62_{\pm.14}$ $0.65_{\pm.15}$ $1.81_{\pm.22}$ $1.85_{\pm.23}$ $0.62_{\pm.13}$ $0.64_{\pm.16}$
CFR $3.15_{\pm.28}$ $3.38_{\pm.30}$ $0.99_{\pm.05}$ $1.01_{\pm.07}$ $\mathbf{1.63_{\pm.21}}$ $\mathbf{1.65_{\pm.25}}$ $0.86_{\pm.18}$ $0.87_{\pm.16}$
GANITE $3.41_{\pm.52}$ $3.67_{\pm.55}$ $0.68_{\pm.13}$ $0.70_{\pm.14}$ $2.48_{\pm.53}$ $2.72_{\pm.59}$ $1.84_{\pm.30}$ $2.02_{\pm.37}$
5.5 Counterfactual Image Generation
Figure 3: PO-Flow counterfactual image generation from female to male faces. $(a)$: Each row shows a factual outcome $y^{(a)}$ (1st column), its encoded latent $z_{y}$ (2nd column), and reverse-time decoding at $t=2/3$, $t=1/3$, and $t=0$ (3rd–5th columns), yielding the counterfactual $\hat{y}_{\text{cf}}^{(1-a)}$. $(b)$: Counterfactuals decoded from perturbed latents $z_{y}+\delta$, with $\delta\sim\mathcal{N}(0,0.2I)$.
Beyond low-dimensional datasets, PO-Flow also succeeds on high-dimensional counterfactual image generation. We train PO-Flow on the latent space of a pre-trained VAE using the CelebA dataset, predicting counterfactual male faces from original female faces. In this setting, $X\in\mathbb{R}^{6}$ encodes visual attributes (hair, eyebrows, bangs, lips), $a\in\{\text{male},\text{female}\}$ denotes gender, and $y\in\mathbb{R}^{512}$ is the VAE latent feature. Visual results are shown in Figures 3 and 6; detailed settings are in Appendix B.2.
5.6 Discussions and Comparisons
Training efficiency: PO-Flow exhibits high training efficiency, converging in approximately 300 gradient steps across multiple causal datasets, as shown in Appendix C.1.
Minimal architecture sensitivity: We find that PO-Flow is the least sensitive to neural network architecture and hyperparameters. Using the same NN structure across all causal datasets, PO-Flow consistently achieves state-of-the-art performance. In contrast, we find that generative approaches including DiffPO, INFs, and GANITE require careful hyperparameter tuning for each dataset.
Unique density estimation: Our continuous normalizing flow (CNF) backbone enables density learning of potential outcomes, as shown in Proposition 4.2 and Section 5.3.1, which, to our knowledge, is the first generative model in the field to achieve this.
Counterfactual: We take a step further by exploring counterfactual outcomes, enabled by the smooth PO-Flow mapping that preserves factual information during encoding. Unlike diffusion-based DiffPO, which has noisy forward processes, PO-Flow offers information-preserving forward processes, as shown in Figure 1. To our knowledge, it is among the first to explore empirical counterfactual prediction without relying on temporal data or explicitly modeling structural causal models.
However, one area for improving counterfactual prediction lies in modeling uncertainty more expressively. The current covariate-conditioned Gaussian perturbations may not fully capture counterfactual variation. Future work could explore more theoretically grounded counterfactual uncertainty.
6 Conclusions
We introduced PO-Flow, a continuous normalizing flow framework for modeling potential and counterfactual outcomes in causal inference. PO-Flow provides a unified generative framework for potential outcome prediction, CATE estimation, and general-purpose counterfactual prediction—without explicitly modeling structural causal models (SCMs) or relying on auxiliary information for the counterfactual task.
Empirically, PO-Flow achieves state-of-the-art performance on benchmark datasets across multiple causal tasks. Among generative models, it is the first to accurately learn the density of potential outcomes, enabling principled uncertainty quantification. Furthermore, PO-Flow is efficient, robust, and extendable to high-dimensional settings such as counterfactual image generation. These results highlight PO-Flow as a unified and scalable solution for causal inference tasks.
Acknowledgment
D.I. acknowledges support from NSF (IIS-2212097), ARL (W911NF-2020-221), and ONR (N00014-23-C-1016). Y.X. acknowledges partial support from NSF DMS-2134037, CMMI-2112533, and the Coca-Cola Foundation. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the sponsor(s).
Appendix A Proofs
Proposition 4.1. Assume that $p_{t}(y|x,a)>0$ for all $y\in\mathbb{R}^{d_{Y}}$ and $t\in[0,1]$. Then, up to a constant independent of $\theta$, the Conditional Flow Matching (CFM) loss and the original Flow Matching (FM) loss are equivalent. Hence,
$\nabla_{\theta}\mathcal{L}_{\text{FM}}(\theta)=\nabla_{\theta}\mathcal{L}_{\text{CFM}}(\theta).$
Proof:
We denote the conditional reference velocity field as $u(y,t,x,a|y_{0})=\frac{d\phi_{t}}{dt}=(1-\sigma_{\text{min}})y_{1}-y_{0}$, as defined in (6). We further define the marginal (unconditional) reference velocity field as:
[…]
where $p(y,t,x,a|y_{0})$ and $p(y,t,x,a)$ denote the probability densities induced by $u(y,t,x,a|y_{0})$ and $u(y,t,x,a)$, respectively. Next, the original flow matching loss, unconditioned on $y_{0}$, is defined as:
$\mathcal{L}_{\text{FM}}=\mathbb{E}_{t\sim\mathcal{U}[0,1],p_{Y_{t},X,A}(\cdot)}\|v_{\theta}(y_{t},t,x,a)-u(y_{t},t,x,a)\|^{2}.$ (15)
In the main text, we chose $\phi_{t}$ as the linear interpolant between $y_{0}$ and $y_{1}$, which results in a constant reference velocity $\frac{d\phi_{t}}{dt}$ independent of $t$. This leads to the simplified CFM loss presented in Eq. (9). However, the general CFM loss, corresponding to an arbitrary reference velocity field $u(y,t\mid y_{0})$ conditioned on $y_{0}$, is defined as:
$\mathcal{L}_{\text{CFM}}=\mathbb{E}_{t\sim\mathcal{U}[0,1],p_{Y}(y_{0}),p_{Y_{t},X,A|Y_{0}}(\cdot|y_{0})}\|v_{\theta}(\cdot)-u_{t}(\cdot|y_{0})\|^{2}$ (16)
Next, we have that:
$\|v_{\theta}(y,t,x,a)-u(y,t,x,a)\|^{2}$ $=\|v_{\theta}(\cdot)\|^{2}-2\langle v_{\theta}(\cdot),u(\cdot)\rangle+\|u(\cdot)\|^{2}$ (17)
$\|v_{\theta}(y,t,x,a)-u(y,t,x,a|y_{0})\|^{2}$ $=\|v_{\theta}(\cdot)\|^{2}-2\langle v_{\theta}(\cdot),u(\cdot|y_{0})\rangle+\|u(\cdot|y_{0})\|^{2}.$
Besides, the term $\mathbb{E}_{t\sim\mathcal{U}[0,1],p_{Y_{t},X,A}(\cdot)}\|v_{\theta}(\cdot)\|^{2}$ can be equivalently expressed in conditional form as:
[…]
$=\int\|v_{\theta}(\cdot)\|^{2}p_{Y_{t},X,A|Y_{0}}(\cdot|y_{0})p_{Y}(y_{0})\,dy_{0}\,dy\,dx\,da\,dt$
$=\mathbb{E}_{t\sim\mathcal{U}[0,1],p_{Y}(y_{0}),p_{Y_{t},X,A|Y_{0}}(\cdot|y_{0})}\|v_{\theta}(\cdot)\|^{2}.$
Furthermore, by the law of total expectation, we have:
[…]
$=\int\left\langle v_{\theta}(\cdot),\int u(\cdot|y_{0})p_{Y_{t},X,A}(\cdot|y_{0})p_{Y}(y_{0})\,dy_{0}\right\rangle\,dy\,dx\,da\,dt$
$=\int\langle v_{\theta}(\cdot),u(\cdot|y_{0})\rangle p_{Y_{t},X,A}(\cdot|y_{0})p_{Y}(y_{0})\,dy_{0}\,dy\,dx\,da\,dt$
$=\mathbb{E}_{t\sim\mathcal{U}[0,1],p_{Y}(y_{0}),p_{Y_{t},X,A|Y_{0}}(\cdot|y_{0})}\langle v_{\theta}(\cdot),u(\cdot|y_{0})\rangle.$
Finally, because $u_{t}(\cdot\mid y_{0})$ does not depend on $\theta$, substituting Eqs. (18) and (19) into Eq. (15) and differentiating with respect to $\theta$ yields:
$\nabla_{\theta}\mathcal{L}_{\text{FM}}(\theta)=\nabla_{\theta}\mathcal{L}_{\text{CFM}}(\theta).$ (20)
Proposition 4.2. Given a noise sample $y_{1}\sim q(\cdot)$ at $t=1$, the log-density of the resulting potential outcome at $t=0$, obtained via PO-Flow, is given by:
$\log\hat{p}_{Y}(\cdot|x,a,y_{1})=\log q(y_{1})+\int_{0}^{1}\nabla\cdot v_{\theta}(y(t),t,x,a)dt.$
Proof:
Samples from Continuous Normalizing Flows (CNFs) evolve according to the following Neural ODE:
$\frac{dy(t)}{dt}=v_{\theta}(y(t),t,x,a),\quad t\in[0,1],$ (21)
which induces a corresponding evolution of the sample density governed by the Liouville continuity equation:
$\partial_{t}p(y,t)+\nabla\cdot\big{(}p(y,t)v_{\theta}(y,t,x,a)\big{)}=0.$ (22)
Here, $p(y,t)$ denotes the time-dependent probability density of the samples $y$.
Next, we have that the dynamics of the density $p(\cdot)$ governed by the velocity field $v_{\theta}(y,t,x,a)$ is given by:
$\frac{d}{dt}\log p(y(t),t)$ $=\frac{\nabla p(y(t),t)\cdot\partial_{t}y(t)+\partial_{t}p(y(t),t)}{p(y(t),t)}$ (23)
$=\frac{\nabla p\cdot\partial_{t}y+\partial_{t}p}{p}\Big{|}_{(y(t),t)}$ (24)
[…]
$=\frac{\nabla p\cdot v_{\theta}-(\nabla p\cdot v_{\theta}+p\nabla\cdot v_{\theta})}{p}\Big{|}_{(y(t),t)}$ (26)
$=-\nabla\cdot v_{\theta}.$ (27)
Starting from an initial sample $y_{1}\sim q(\cdot)$ and integrating from $t=1$ to $t=0$, we have:
$\log\hat{p}_{Y}(\cdot|x,a,y_{1})=\log q(y_{1})+\int_{0}^{1}\nabla\cdot v_{\theta}(y(t),t,x,a)dt.$ (28)
Besides, if $q(\cdot)$ is selected as the commonly used $N(0,I)$, we have:
$\log\hat{p}_{Y}(\cdot|x,a,y_{1})=-\frac{d_{Y}}{2}\log 2\pi-\frac{1}{2}\|y_{1}\|^{2}+\int_{0}^{1}\nabla\cdot v_{\theta}(y(t),t,x,a)dt.$ (29)
Appendix B Experimental Details
B.1 Synthetic Causal Datasets
Counterfactual outcomes are unobservable in real-world data but can be simulated using synthetic datasets. In our experiments, we evaluate on widely used benchmarks, including the Atlantic Causal Inference Conference datasets (ACIC 2016 and ACIC 2018), the Infant Health and Development Program (IHDP) [24, 42], and the IBM Causal Inference Benchmark [56].
The ACIC 2016 dataset contains 4,802 samples with 82 covariates each ($d_{X}=82$), ACIC 2018 has 1,000 samples with 177 covariates, IHDP includes 747 samples with 25 covariates, and the IBM dataset consists of 1,000 samples with 177 covariates. Each dataset provides three synthetic values per individual: the factual outcome $y^{(a)}$, the assigned treatment $a$, and the mean potential outcomes under both treatments, $\mu_{0}$ (for $a=0$) and $\mu_{1}$ (for $a=1$).
Moreover, under the following dataset synthesis assumptions: (1) fully specified outcome model $Y^{(A)}=f(X,A)+\epsilon$; (2) shared noise across treatment arms for each individual; and (3) invertibility of the structural function such that, given $X$, $A$, and $Y^{(A)}$, one can solve for $\epsilon$ — we can synthesize the counterfactual outcome through the following procedure:
1. Abduction: infer the unobserved noise term $\epsilon^{\prime}$ from the factual outcome $Y^{(A)}$, covariates $X$, and treatment $A$;
2. Action: intervene by changing the treatment assignment to $1-A$;
3. Prediction: use the inferred noise $\epsilon^{\prime}$ and new treatment to compute the counterfactual outcome via $Y^{(1-A)}=f(X,1-A)+\epsilon^{\prime}$.
All datasets are generated using diverse simulation mechanisms, and we provide the specific versions used along with our code. For example, the IHDP dataset is generated using 25 covariates and a coefficient matrix $W$, following the data-generating process:
$X\sim\text{Real-world distribution }p_{X}(\cdot),$ (30)
$A\sim\text{Real-world Treatment assignment based on}\ X,$
$Y=A(X\beta-\omega)+(1-A)\exp((X+W)\beta)+\epsilon,\quad\epsilon\sim N(0,1).$
where $\beta$, $W$, and $\omega$ are parameters defined by the specific simulation setting.
B.2 Counterfactual Image Experiments
To intuitively visualize PO-Flow’s counterfactual performance and evaluate its effectiveness on high-dimensional data, we train PO-Flow on the latent space of a pre-trained Variational Autoencoder (VAE) [36] using the CelebA dataset.
The CelebFaces Attributes Dataset (CelebA) comprises 202,599 face images, each with a resolution of $178\times 218$ pixels. Each image is annotated with 40 binary facial attributes. In our experiments, we selected the male/female attribute as the treatment variable $A\in\{\text{male},\text{female}\}$, and defined covariates $X\in\mathbb{R}^{6}$ to represent features such as hair, eyebrows, bangs, and lips.
The VAE is pre-trained on the full CelebA dataset with a 512-dimensional latent space. Each $178\times 218$ face image is then encoded into this latent space, yielding the training target $Y\in\mathbb{R}^{512}$. PO-Flow is subsequently trained on the triplet $(Y,X,A)$ to generate counterfactual male face representations given original female faces.
B.3 Neural Network Structure and Implementation Information
The neural network used in experiments on causal datasets (except the CelebA image task) takes as input $y(t)$, $x$, $t$, and $a$. The model is composed of the following components: (1) an outcome embedding layer that projects $y(t)$ to the same dimension as the concatenated $(x,a)$; (2) a FiLM layer that applies feature-wise affine transformations to the embedded $y(t)$, conditioned on $(x,a)$; (3) two residual blocks, each consisting of a gated two-layer MLP conditioned on $(x,a)$, with residual averaging and skip connections; and (4) a projection layer that maps the final hidden representation to two velocity vectors $v_{0}$ and $v_{1}$, corresponding to treatments 0 and 1.
The neural network used for the counterfactual image generation task has an architecture composed of the following components: (1) a conditioning MLP that encodes the pair $(t,a)$ into a hidden feature vector via two linear layers and a ReLU activation; (2) an input layer that takes the concatenation of $y(t)$ and $(t,a)$, followed by a ReLU-activated linear transformation; (3) four residual blocks, each composed of two linear layers with a ReLU nonlinearity, where the conditioning feature is injected additively at each block; and (4) an output linear layer that maps the final hidden representation back to the latent outcome dimension.
All experiments are conducted on a V100 GPU with 16 GB of memory. The training time for causal datasets (excluding CelebA counterfactual face image generation) depends on the dataset and batch sizes. For example, training on the ACIC 2018 dataset with a batch size of 256 typically takes 2–4 seconds per epoch. Log-density evaluation, as discussed in Section 4.3, involves divergence integration and requires significantly more time than other test metrics.
B.4 Runge-Kutta numerical integration method
We apply the classical 4th-order Runge-Kutta method to solve the Neural ODE in the forward process of PO-Flow. Given a factual outcome $y^{(a)}$, we aim to compute the latent embedding $z_{y}=y(1)$ by integrating the velocity field $v_{\theta}$ from $t=0$ to $t=1$, with initial condition $y(0)=y^{(a)}$. Let $h$ denote the step size, and define discrete time steps $t_{k}=kh$, for $k=0,1,\dots,T$, such that $Th=1$. The update rule at each step is:
$\small y(t_{k+1})=y(t_{k})+\frac{h}{6}\left(k_{1}+2k_{2}+2k_{3}+k_{4}\right),$
where
$k_{1}$ $=v_{\theta}(y(t_{k}),t_{k},x,a),\ k_{2}=v_{\theta}\left(y(t_{k})+\frac{h}{2}k_{1},t_{k}+\frac{h}{2},x,a\right),$
$k_{3}$ $=v_{\theta}\left(y(t_{k})+\frac{h}{2}k_{2},t_{k}+\frac{h}{2},x,a\right),\ k_{4}=v_{\theta}\left(y(t_{k})+hk_{3},t_{k}+h,x,a\right),$
with $y(0)=y^{(a)}$ and $y(1)=z_{y}$. This produces a numerically integrated trajectory from the observed sample to its latent representation.
B.5 Hutchinson trace estimator
The computations of the log-density of the POs involve the calculation of $\nabla\cdot v_{k}(x,t)$, i.e., the divergence of the velocity field represented by a neural network. This may be computed by brute force using reverse-mode automatic differentiation, which is much slower and less stable in high dimensions.
We can express $\nabla\cdot v_{\theta}(y,t,x,a)=\mathbb{E}_{\epsilon\sim p(\epsilon)}\left[\epsilon^{T}J_{v}(y)\epsilon\right]$, where $J_{v}(y)$ is the Jacobian of $v_{\theta}(y,t,x,a)$ at $y$. Given a fixed $\epsilon$, we have $J_{v}(y)\epsilon=\lim_{\sigma\to 0}\frac{v_{\theta}(y+\sigma\epsilon,t,x,a)-v_{\theta}(y,t,x,a)}{\sigma}$, which is the directional derivative of $v_{\theta}$ along the direction $\epsilon$. Therefore, for a sufficiently small $\sigma>0$, we can propose the following Hutchinson estimator  [31, 64]:
$\small\nabla\cdot v_{\theta}(y(t),t,x,a)\approx\mathbb{E}_{p(\epsilon)}\left[\epsilon\cdot\frac{v_{\theta}(y(t)+\sigma\epsilon,t,x,a)-v_{\theta}(y(t),t,x,a)}{\sigma}\right],$ (31)
where $p(\epsilon)$ is a distribution in $\mathbb{R}^{d_{Y}}$ satisfying $\mathbb{E}[\epsilon]=0$ and Cov$(\epsilon)=I$ (e.g., a standard Gaussian). This approximation becomes exact as $\sigma\to 0$.
B.6 Adapted Implementation Details for DiffPO and GANITE Counterfactuals
DiffPO was not originally designed for counterfactual tasks. However, since DiffPO uses a diffusion model as its base, we are able to adapt the algorithm to incorporate an "encoding" and "decoding" structure similar to that of PO-Flow. Our adapted implementation of their algorithm for counterfactual estimation is summarized in Algorithm 3.
  Input: $(y^{(a)},x,a)\sim p_{Y,X,A}(\cdot)$
  1: $z_{0}\leftarrow y^{(a)}$
  2: Forward Process:
  3:     for $t=0,\cdots,T-1$:
  4:       $z_{t+1}\leftarrow\sqrt{\frac{\alpha_{t+1}}{\alpha_{t}}}z_{t}+f_{\theta}(z_{t},t|x,a)\left(\sqrt{1-\alpha_{t+1}}-\sqrt{\frac{\alpha_{t+1}(1-\alpha_{t})}{\alpha_{t}}}\right)$
  5:     $z_{y}\leftarrow z_{T}$
  6: Reverse Process:
  7:     Follow their official implementation in Section 5.3
  8: Return $\hat{y}^{(1-a)}:=\hat{y}_{0}$
Algorithm 3 Adapted DiffPO for Counterfactual
Similarly, GANITE was not originally designed for counterfactual inference. It consists of a generator and a discriminator, with the final output targeting CATE estimation. However, the generator also produces a proxy counterfactual given the factual input, which we use to evaluate its counterfactual estimation performance.
Appendix C Additional Experimental Results
C.1 RMSE Convergence
PO-Flow demonstrates rapid convergence across datasets, requiring only around 250 training iterations (not epochs) with a batch size of 200 per iteration. This highlights the model’s strong training efficiency.
C.2 Additional counterfactual images
Figure 6: PO-Flow counterfactual image generation from female to male faces. $(a)$: Each row shows a factual outcome $y^{(a)}$ (1st column), its encoded latent $z_{y}$ (2nd column), and reverse-time decoding at $t=2/3$, $t=1/3$, and $t=0$ (3rd–5th columns), yielding the counterfactual $\hat{y}^{(1-a)}$. $(b)$: Counterfactuals decoded from perturbed latents $z_{y}+\delta$, with $\delta\sim\mathcal{N}(0,0.2I)$.